Find an equation for the line perpendicular to the line −2x + 2y = −7 having the same y-intercept as −7x + 9y = 63
Answer by Alan3354(69443)

Find an equation for the line perpendicular to the line −2x + 2y = −7 having the same y-intercept as −7x + 9y = 63
----------
Step 1, find the slope of -2x + 2y = -7
The slope of lines perpendicular is the negative inverse.
----------
Step 2, find the y-int of -7x + 9y = 63
-----
Step 3, y = mx + b
========================
It would be to your advantage to know how to do this.
I don't need the practice.